OpenStudy (anonymous):

If you shake one end of a taut rope steadily six times each second, what would be the period of the sinusoidal wave set up in the rope?

OpenStudy (shamim):

So frequency f=6 Hz

OpenStudy (shamim):

We know time period T=1/f=?

OpenStudy (shamim):

Response plz

OpenStudy (anonymous):

its ms that im looking for

OpenStudy (shamim):

T=1/6=?

OpenStudy (shamim):

ms means?

OpenStudy (anonymous):

i tired 1/6 and it said it was wrong

OpenStudy (shamim):

1/6=0.16667